### v1

@@ Abstract
Large Language Models (LLMs) have shown considerable potential in automating decision logic within knowledge-intensive processes. However, their effectiveness largely depends on the strategy and quality of prompting. Since decision logic is typically embedded in prompts, it becomes challenging for end users to modify or refine it. Decision Model and Notation (DMN) offers a standardized graphical approach for defining decision logic in a structured, user-friendly manner. This paper introduces a DMN-guided prompting framework that breaks down complex decision logic into smaller, manageable components, guiding LLMs through structured decision pathways. We implemented the framework in a graduate-level course where students submitted assignments. The assignments and DMN models representing feedback instructions served as inputs to our framework. The instructor evaluated the generated feedback and labeled it for performance assessment. Our approach demonstrated promising results, outperforming chain-of-thought (CoT) prompting. Students also responded positively to the generated feedback, reporting high levels of perceived usefulness in a survey based on the Technology Acceptance Model.
@@
Large Language Models, Decision Model and Notation, Prompt Engineering, Automated Feedback
@@ 1 Introduction
Large Language Models (LLMs) have shown considerable potential in automating decision logic within knowledge-intensive processes [12]. Their capabilities are increasingly leveraged across multiple domains where complex decisions must be made. However, the effectiveness of LLMs critically depends on the quality of their instructions, a.k.a, prompts. Developing such prompts—commonly known as prompt engineering—is often a trial-and-error process, requiring technical expertise and posing challenges related to transparency, reusability, and maintenance [15, 19]. As decision logics evolve, their instructional prompts must also be adapted—a task often non-trivial for business users lacking technical proficiency in developing or adjusting prompts.
To address these challenges, we propose a novel prompting strategy that integrates Decision Model and Notation (DMN) into prompt definitions. DMN is a standardized graphical notation for modeling decision logic [5, 1]. By leveraging DMN to modularize and guide prompt structures, our strategy enables LLMs to follow structured decision pathways, breaking down complex logic into interpretable and manageable components. This approach reduces the dependency on ad hoc prompt engineering, making it easier to understand, modify, and maintain—particularly in settings where clarity, consistency, and collaboration are essential.
We operationalized this idea in a DMN-guided prompting framework configured in a graduate-level course to automatically support generating feedback for students’ assignments while having the instructor in the loop. Students submitted graphical models representing domain-specific decisions. These models were transformed into natural texts, which are given to our framework in addition to instructional DMN rules defined by the course instructor. The framework generated feedback, which was given to students after the instructor labeled them so students could distinguish between correct and incorrect feedback. These labeled data were used to evaluate the approach, and the results show that DMN-guided prompting outperformed equivalent prompts defined purely using chain-of-thought prompting [20]. In addition, students reported high levels of perceived usefulness, as measured through a survey designed based on the Technology Acceptance Model (TAM) [4].
@@ 2 Background
This section briefly details the “Decision Model and Notation” and “Prompt Engineering”.
@@ 2.1 Decision Model and Notation
Figure 1 presents a sample DRD that will serve as a running example throughout this paper to illustrate our approach (for the complete DMN specification, refer to [1]). This diagram specifies the rules for a fictitious loan eligibility process, where the decision is based on the applicant’s Salary and Credit Score, modeled as input data elements. A decision table element, named Loan Approval Rule, encodes the decision logic. The decision output is then passed to a literal expression element, called Loan Approval Result. Literal expressions represent decision logic using textual statements—in this case, defining the message to be displayed to the user based on the decision outcome. Loan Approval is an annotation that does not have any semantics yet enables providing more information about elements in the DMN model.
@@ 2.2 Prompt Engineering
Among these strategies, Chain-of-thought prompting is reported to be effective for complex tasks as it follows the divide-and-conquer approach, decomposing a task into smaller sub-tasks and helping LLMs to follow a step-by-step approach to solve a problem. For example, instead of prompting the LLM to evaluate an assignment holistically, the task can be divided into sub-prompts that independently assess correctness, clarity, and completeness. These individual judgments are combined to form the final feedback [3]. The divide-and-conquer approach is particularly relevant in domains where structured reasoning and consistency are important, such as educational assessment or business rule evaluation. However, applying this method manually, similar to Chain-of-Thought, can be challenging. It often requires writing and managing multiple prompts and coordinating their output coherently.
@@ 3 DMN-Guided Prompting
Although DMN supports multiple representations of decision logic, our approach adopts a divide-and-conquer strategy—including chain-of-thought prompting [20]—by organizing logic into modular units we refer to as triples. Each triple comprises (1) a set of input data elements, (2) a decision table, and (3) a corresponding literal expression. A simple illustration in Figure 1 includes one triple. These triples encapsulate logically self-contained decision-making units, each functioning as an independent sub-prompt. This modular structure enables LLMs to process each unit sequentially and coherently.
@@
Figure 4 outlines the full prompt used to parse and evaluate the running example DMN file. The process is divided into four sequential parts. While the complete contents of Part B and Part C are not shown in the figure, the roles of each part are as follows:
@@ 3.3 Result Compilation and Final Output
After completing all sub-tasks, the LLM assembles the outputs into a single response as described in Part D. Each triple is evaluated independently, and the results are synthesized into a coherent final output. The full implementation, including the Python source code, the running example DMN file, and example input texts, is available on our GitHub repository 111 https://github.com/Shaghayegh-Abedi/DMN-prompting-project. It is important to emphasize that there is no need to modify the prompt when changing the decision logic. Instead, users modify the DMN model, which the LLM interprets to evaluate the logic and generate the results.
@@ 4 Case Study
This case study investigates the feasibility and effectiveness of the proposed DMN-guided prompting framework for automated feedback generation in an academic context. Specifically, it examines how decision models can structure the behavior of large language models (LLMs) to evaluate student-created process models and return predefined feedback addressing common known problems in the models. We begin by describing the educational setting, followed by the implementation and evaluation methodology, and conclude with findings concerning system performance and student perceptions.
@@ 4.1 Educational Setting and Study Context
4.2 Framework Deployment in the Case Study
Figure 7 depicts deploying the DMN-guided prompting framework within the course. The setup comprises three primary actors: (i) the instructor, who encodes feedback logic using DMN triples; (ii) the student, who submits process models for evaluation; and (iii) the DMN-guided prompting system, which generates feedback based on the encoded logic for each submission.
To ensure reliability, the generated feedback was not directly shared with students. Instead, a human-in-the-loop mechanism was implemented: instructors manually reviewed each comment and classified it as correct or incorrect. The labeled feedback was subsequently communicated to students. This approach safeguarded the instructional quality and enabled an empirical evaluation of the system’s predictive accuracy. The framework was executed for robustness using two LLMs—GPT-4o (via the OpenAI API) and Gemini 1.5 Pro (via Google’s Gemini API). Additionally, we implemented a Chain-of-Thought (CoT) [20] prompting set up as a performance baseline, allowing us to isolate the impact of the DMN structure from general prompt engineering strategies.
@@ 4.3 Technical Implementation
We developed a software component to operationalize the framework that translates Petri net models into structured natural language descriptions. These descriptions retain critical control flow and resource semantics, enabling LLMs to interpret process models as standardized textual artifacts. The implementation is publicly available on GitHub for reuse in similar applications 222 https://github.com/Shaghayegh-Abedi/DMN-prompting-project. The core feedback mechanism comprises nine decision rules, each encoded as a table in the Decision Model and Notation (DMN). These rules align with specific BPR principles and provide pedagogically meaningful evaluation criteria for student submissions.
Figure 8: Overview of instructor-labeled feedback per group and rule.
@@ 4.4 Evaluation Results
Figure 8 presents an overview of instructor-labeled feedback (for GPT-4o experiment) across student groups and nine BPR rules. Each rule is assessed along two dimensions: whether the student correctly applied the principle (Rule column), and whether the LLM-generated feedback accurately reflected this (Feedback column). Table 1 provides a comparative analysis of four configurations: DMN-guided and CoT prompting strategies, each paired with GPT-4o and Gemini 1.5 Pro. Standard classification metrics—Precision, Recall, F1-score, and Accuracy—were computed by comparing system outputs with instructor labels.
The specified results in Table 1) show that the DMN-guided prompting consistently outperforms the CoT baseline across all metrics and models. The GPT-4o + DMN configuration yielded the highest performance: 0.91 precision, 0.90 recall, 0.91 F1-score, and 0.87 accuracy. In contrast, GPT-4o + CoT achieved perfect recall (1.00) but suffered a severe drop in precision (0.36), resulting in a markedly lower F1-score of 0.53—suggesting a high rate of false positives.
@@ 4.5 Feedback Analysis and Rule-Specific Insights
To complement the quantitative findings, we conducted a detailed qualitative analysis to assess the system’s capability to detect individual BPR principles and reveal underlying student comprehension patterns.
All 14 groups correctly applied Rule 1 (triage), and the DMN-guided framework using GPT-4o flawlessly recognized these instances.
@@
Rule 2 (task composition) was correctly implemented by the majority (79%) of the groups, and the framework achieved 86% accuracy in identifying them—closely aligned with the overall accuracy. Notably, in the two misclassified cases, students labeled the merged task using unexpected phrasing (e.g., “check and register result” instead of the expected formulation indicating the task is related to credit check), leading to reduced model confidence in mapping the task semantics, as illustrated in Fig. 9.
For Rule 3 (knock-out), 93% of the groups modeled the principle correctly, and the framework matched this performance with a 93% accuracy rate—exceeding its overall average. This rule entails rejecting cases early based on credit score assessments (see the XOR-split after “Check and Register Result” in Fig. 9).
@@
Regarding Rule 7 (parallelism between two automated tasks), 71% of the groups applied it correctly, and the framework reached 79% accuracy. Some confusion arose when students incorrectly applied task composition—merging two external web services into a single task- violating design control assumptions, as the process designer cannot merge such services.
Finally, for Rules 8 (parallelism of an automated and a manual task) and 9 (sequence within parallelized tasks), 86% of the groups applied them correctly, and the framework achieved an identical 86% accuracy—matching its overall performance.
@@ 4.6.1 Note on Perceived Ease of Use:
Perceived Ease of Use (PEU) was not assessed in this study, as students interacted with the system indirectly via the university’s Learning Management System (LMS). They neither directly accessed nor operated the system interface, rendering PEU evaluation inapplicable.
@@ 5 Conclusion
This paper presented a novel prompting framework that integrates Decision Model and Notation (DMN) with Large Language Models (LLMs) to guide decision logic in a modular, interpretable, and user-friendly manner. By structuring prompts as DMN triples—each consisting of input data, decision tables, and literal expressions—our approach operationalizes a divide-and-conquer strategy for LLM behavior control. The case study conducted in a graduate-level course demonstrated that DMN-guided prompting consistently outperformed chain-of-thought (CoT) methods in predictive accuracy. Regarding user-perceived usefulness, students reported high perceived value from the feedback generated by our DMN-based approach.
The results confirm the viability of DMN as a low-code interface for LLM control in educational settings, particularly for structured feedback generation. More broadly, this work illustrates the untapped potential of DMN to encode domain-specific reasoning in a format that LLMs can interpret and execute reliably, offering a promising direction for rule-based, transparent AI systems.
Future research can expand in several directions. First, exploring how DMN-guided prompting can be adapted for other application domains, such as healthcare or compliance auditing, will help assess its generalizability. Second, combining DMN with retrieval-augmented generation (RAG) techniques may further enhance reasoning grounded in external knowledge. Third, if customized message generation is desired, the LLM can post-process the output messages using a secondary prompt; otherwise, the framework returns only the predefined messages specified in the DMN logic. Fourth, this approach can be complemented by incorporating few-shot learning, where example inputs are used to improve LLM accuracy when interpreting or applying DMN rules. Fifth, other ways of defining DMN rules and alternative strategies for prompt construction can be explored to increase flexibility and expressiveness. Finally, investigating multi-turn interactions and conversational agents powered by DMN-guided LLMs could open new opportunities for explainable and controllable AI interfaces in decision support systems.

### v2

@@ Abstract
Large Language Models (LLMs) have shown considerable potential in automating decision logic within knowledge-intensive processes. However, their effectiveness largely depends on the strategy and quality of prompting. Since decision logic is typically embedded in prompts, it becomes challenging for end users to modify or refine it. Decision Model and Notation (DMN) offers a standardized graphical approach for defining decision logic in a structured, user-friendly manner. This paper introduces a DMN-guided prompting framework that breaks down complex decision logic into smaller, manageable components, guiding LLMs through structured decision pathways. We implemented the framework in a graduate-level course where students submitted assignments. The assignments and DMN models representing feedback instructions served as inputs to our framework. The instructor evaluated the generated feedback and labeled it for performance assessment. Our approach demonstrated promising results, outperforming chain-of-thought (CoT) prompting in our case study. Students also responded positively to the generated feedback, reporting high levels of perceived usefulness in a survey based on the Technology Acceptance Model.
@@
Large Language Models, Decision Model and Notation, Automated Feedback, Prompt Engineering
@@ 1 Introduction
Large Language Models (LLMs) have shown considerable potential in automating decision logic within knowledge-intensive processes [12]. Their capabilities are increasingly leveraged across multiple domains where complex decisions must be made. However, the effectiveness of LLMs critically depends on the quality of their instructions, also known as prompts. Developing such prompts (commonly known as prompt engineering) is often a trial-and-error process, requiring technical expertise and posing challenges related to transparency, reusability, and maintenance [15, 19]. As decision logics evolve, their instructional prompts must also be adapted - a task often non-trivial for business users lacking technical proficiency in developing or adjusting prompts.
To address these challenges, we propose a novel framework that integrates Decision Model and Notation (DMN) into prompt template definitions. DMN is a standardized graphical notation for modeling decision logic [5, 1]. By leveraging DMN to modularize and guide prompt structures, our strategy enables LLMs to follow structured decision pathways, breaking down complex logic into interpretable and manageable components. This approach reduces the dependency on ad hoc changes of prompts, making it easier to understand, modify, and maintain - particularly in settings where clarity, consistency, and collaboration are essential.
We implemented our proposed framework as a DMN-guided prompting framework configured in a graduate-level course to automatically support generating feedback for students’ assignments while having the instructor in the loop. Students submitted graphical models representing domain-specific decisions. These models were transformed into natural texts, which are given to our framework in addition to instructional DMN rules defined by the course instructor. The framework generated feedback, which was given to students after the instructor labeled them, so students could distinguish between correct and incorrect feedback. These labeled data were used to evaluate the approach, and the results show that DMN-guided prompting outperformed equivalent prompts defined purely using chain-of-thought prompting [20]. In addition, students reported high levels of perceived usefulness, as measured through a survey designed based on the Technology Acceptance Model (TAM) [4].
@@ 2.1 Decision Model and Notation
Figure 1 presents a sample DRD that will serve as a running example throughout this paper to illustrate our approach (for the complete DMN specification, refer to [1]). This diagram specifies the rules for a fictitious loan eligibility process, where the decision is based on the applicant’s Salary and Credit Score, modeled as input data elements. A decision table element, named Loan Approval Rule, encodes the decision logic. The decision output is then passed to a literal expression element, called Loan Approval Result. Literal expressions represent decision logic using textual statements - in this case, defining the message to be displayed to the user based on the decision outcome. Loan Approval is an annotation that does not have any semantics yet enables providing more information about elements in the DMN model.
@@ 2.2 Prompt Engineering
Among these strategies, Chain-of-thought prompting is reported to be effective for complex tasks as it follows the divide-and-conquer approach, decomposing a task into smaller sub-tasks and helping LLMs to follow a step-by-step approach to solve a problem. For example, instead of prompting the LLM to evaluate an assignment holistically, the task can be divided into sub-prompts that independently assess correctness, clarity, and completeness. These individual judgments are combined to form the final feedback [3]. Dividing prompts into individual parts is effective, yet some parts, like business logics, change more frequently than others, requiring the prompt engineer to refine the prompts. A framework enabling the separation of such business rules using DMN can help business people to apply the changes directly, requiring less intervention from a prompt engineer, and enabling organizations to apply changes more easily and quickly.
@@ 3 DMN-Guided Prompting
Although DMN supports multiple representations of decision logic, our approach adopts a divide-and-conquer strategy (including chain-of-thought prompting [20]) by organizing logic into modular units we refer to as triples. Each triple comprises (1) a set of input data elements, (2) a decision table, and (3) a corresponding literal expression (as illustrated early in the paper in Figure 1, including one sample triple). These triples encapsulate logically self-contained decision-making units, each functioning as an independent sub-prompt. This modular structure enables LLMs to process each unit sequentially and coherently.
@@
Figure 4 outlines the full prompt template used to parse and evaluate the running example DMN file, which can be adopted for different use cases differently. Prompt templates refinement is important as it gives organizations freedom in how to design DMN models based on their needs. The process is divided into four sequential parts. While the complete contents of Part B and Part C are not shown in the figure, the roles of each part are as follows:
@@ 3.3 Result Compilation and Final Output
After completing all sub-tasks, the LLM assembles the outputs into a single response as described in Part D. Each triple is evaluated independently, and the results are synthesized into a coherent final output. The full implementation, including the Python source code, the running example DMN file, and example input texts, is available on our GitHub repository 111 https://github.com/Shaghayegh-Abedi/DMN-prompting-project. It is important to emphasize that there is no need to modify the prompt if decision rules are changed. Instead, users modify the DMN model, which the LLM interprets to evaluate the logic and generate the results.
@@ 4 Case Study
This case study investigates the feasibility and effectiveness of the proposed DMN-guided prompting framework for automated feedback generation in an academic context. Specifically, it examines how decision models can structure the behavior of large language models (LLMs) to evaluate student-created process models and return predefined feedback addressing common known problems in the models. We begin by describing the educational setting, followed by the implementation and evaluation methodology, and conclude with findings concerning system performance and student perceptions. Due to privacy concerns, the student submissions and responses could not be published. However, the framework and example materials are provided in our GitHub repository to help readers apply the framework to their own use cases.
@@ 4.1 Educational Setting and Study Context
4.2 Framework Deployment in the Case Study
Figure 7 depicts deploying the DMN-guided prompting framework within the course. The setup comprises four primary actors: (i) the prompt engineer defining the prompt template, (ii) the instructor, who encodes feedback logic using DMN triples; (iii) the student, who submits process models for evaluation; and (iv) the DMN-guided prompting system, which generates feedback based on the encoded logic for each submission.
@@ 4.2 Framework Deployment in the Case Study
To ensure reliability, the generated feedback was not directly shared with students. Instead, a human-in-the-loop mechanism was implemented: instructors manually reviewed each comment and classified it as correct or incorrect. The labeled feedback was subsequently communicated to students, reducing the risk of misleading them into assuming that the received feedback was always correct. This approach safeguarded the instructional quality and enabled an empirical evaluation of the system’s predictive accuracy. The framework was executed for robustness using two LLMs - GPT-4o (via the OpenAI API) and Gemini 1.5 Pro (via Google’s Gemini API). Additionally, we implemented a Chain-of-Thought (CoT) [20] prompting set up as a performance baseline, allowing us to isolate the impact of the DMN structure from general prompt engineering strategies.
@@ 4.3 Technical Implementation
We developed a software component to operationalize the framework that translates Petri net models into structured natural language descriptions. These descriptions retain critical control flow and resource semantics, enabling LLMs to interpret process models as standardized textual artifacts. The implementation is publicly available on GitHub for reuse in similar applications. The core feedback mechanism comprises nine decision rules, each encoded as a table in the Decision Model and Notation (DMN). These rules align with specific BPR principles and provide pedagogically meaningful evaluation criteria for student submissions.
Figure 8: Overview of instructor-labeled evaluations of GPT-4o–generated feedback, presented across student groups and nine BPR rules. Each rule is assessed along two dimensions: whether the student correctly applied the principle (Rule column) and the GPT-4o feedback (Feedback column).
@@ 4.4 Evaluation Results
Figure 8 presents an overview of instructor-labeled feedback (for GPT-4o experiment) across student groups and nine BPR rules. Each rule is assessed along two dimensions: the student correctly applied the principle (Rule column), and the LLM-generated feedback (Feedback column). Table 1 provides a comparative analysis of four configurations: DMN-guided and CoT prompting strategies, each paired with GPT-4o and Gemini 1.5 Pro. Standard classification metrics - Precision, Recall, F1-score, and Accuracy - were computed by comparing system outputs with instructor labels.
The specified results in Table 1 show that the DMN-guided prompting generally outperforms the CoT baseline across most metrics and models. The GPT-4o + DMN configuration yielded the highest performance: 0.91 precision, 0.90 recall, 0.91 F1-score, and 0.87 accuracy. In contrast, GPT-4o + CoT achieved perfect recall (1.00) but suffered a severe drop in precision (0.36), resulting in a markedly lower F1-score of 0.53 - suggesting a high rate of false positives.
@@ 4.5 Feedback Analysis and Rule-Specific Insights
To complement the quantitative findings, we conducted a detailed qualitative analysis to assess the system’s capability to detect individual BPR principles and reveal underlying student comprehension patterns. All 14 groups correctly applied Rule 1 (triage), and the DMN-guided framework using GPT-4o flawlessly recognized these instances.
@@
Rule 2 (task composition) was correctly implemented by the majority (79%) of the groups, and the framework achieved 86% accuracy in identifying them - closely aligned with the overall accuracy. Notably, in the two misclassified cases, students labeled the merged task using unexpected phrasing (e.g., “check and register result” instead of the expected formulation indicating the task is related to credit check), leading to reduced model confidence in mapping the task semantics, as illustrated in Fig. 9.
For Rule 3 (knock-out), 93% of the groups modeled the principle correctly, and the framework matched this performance with a 93% accuracy rate - exceeding its overall average. This rule entails rejecting cases early based on credit score assessments (see the XOR-split after “Check and Register Result” in Fig. 9).
@@
Regarding Rule 7 (parallelism between two automated tasks), 71% of the groups applied it correctly, and the framework reached 79% accuracy. Some confusion arose when students incorrectly applied task composition - merging two external web services into a single task- violating design control assumptions, as the process designer cannot merge such services.
Finally, for Rules 8 (parallelism of an automated and a manual task) and 9 (sequence within parallelized tasks), 86% of the groups applied them correctly, and the framework achieved an identical 86% accuracy - matching its overall performance.
@@ 4.6.1 Note on Perceived Ease of Use:
Perceived Ease of Use (PEU) was not assessed, since students interacted with the system only through the university’s Learning Management System (LMS). In this context, PEU would reflect the usability of the LMS rather than our system, making it irrelevant. In contrast, Perceived Usefulness (PU) was evaluated, as students could reflect on the usefulness of the feedback received.
4.6.2 Threats to Validity
Our study has several limitations. Internal validity may be affected by the manual labeling of feedback, which introduces subjectivity despite the human-in-the-loop review. Construct validity is constrained by the chosen metrics (precision, recall, F1, accuracy) and the survey-based perceived usefulness, which may not fully capture long-term learning outcomes. External validity is limited since the case study was conducted in a single graduate-level course with 24 groups and nine decision rules, reducing generalizability to other domains and contexts. Finally, conclusion validity may be influenced by the relatively small dataset and ceiling effects for some rules. Further studies in diverse educational and professional settings are needed to strengthen the generalizability of our findings.
@@ 5 Conclusion
This paper presented a novel prompting framework that integrates Decision Model and Notation (DMN) with Large Language Models (LLMs) to guide decision logic in a modular, interpretable, and user-friendly manner. By structuring prompts as DMN triples (each consisting of input data, decision tables, and literal expressions), our approach operationalizes a divide-and-conquer strategy for LLM behavior control. The case study conducted in a graduate-level course demonstrated that DMN-guided prompting generally outperformed chain-of-thought (CoT) methods in predictive accuracy. Regarding user-perceived usefulness, students reported high perceived value from the feedback generated by our DMN-based approach.
The results confirm the viability of DMN for LLM control in educational settings, particularly for structured feedback generation. More broadly, this work illustrates the untapped potential of DMN to encode domain-specific reasoning in a format that LLMs can interpret and execute reliably, offering a promising direction for rule-based, transparent AI systems.
Future research can expand in several directions. First, exploring how DMN-guided prompting can be adapted for other application domains, such as healthcare or compliance auditing, will help assess its generalizability. Second, combining DMN with retrieval-augmented generation (RAG) may further enhance reasoning grounded in external knowledge. Third, if customized message generation is desired, the LLM can post-process the output messages using a secondary prompt; otherwise, the framework returns only the predefined messages specified in the DMN logic. Fourth, this approach can be complemented by incorporating few-shot learning, where example inputs are used to improve LLM accuracy when interpreting or applying DMN rules. Fifth, other ways of defining DMN rules and alternative strategies for prompt construction can be explored to increase flexibility and expressiveness. Finally, investigating multi-turn interactions and conversational agents powered by DMN-guided LLMs could open new opportunities for explainable and controllable AI interfaces in decision support systems.
5.0.1 Acknowledgment:
The computations for testing and developing our approach were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Appendix 0.A Appendix
This section provides a list of questions asked in the survey to measure perceived usefulness, following the Technology Acceptance Model (TAM). Participants rated their agreement with the following statements on a 7-point Likert scale (1 = Strongly Disagree, 7 = Strongly Agree):
  1. 1.
    Using the auto-feedback system enabled me to identify relevant BPR patterns more quickly.
  2. 2.
    Using the auto-feedback system improved my performance in applying BPR patterns to business process redesign.
  3. 3.
    Using the auto-feedback system increased my productivity when analyzing and revising business process solutions.
  4. 4.
    Using the auto-feedback system enhanced my effectiveness in recognizing missing BPR patterns.
  5. 5.
    Using the auto-feedback system made it easier for me to understand which BPR patterns were applicable.
  6. 6.
    I found the auto-feedback system useful for learning and applying BPR patterns in business process redesign tasks.